# Improving Generalization of Neural Vehicle Routing Problem Solvers Through the Lens of Model Architecture

### v2

@@ 5.2 Effectiveness Analysis of the DS Decoder
5.3 Effectiveness Analysis of ESF + DS
To further evaluate the performance of our proposed ESF and DS decoder in real-world scenarios, we choose the widely recognized TSPLIB [11] and CVRPLIB [12] as the benchmarking datasets for testing. Considering the training sizes of AMDKD ($n_{\textit{tr}}\in\{50,100\}$) and OMNI-VRP ($n_{\textit{tr}}\in[50,200]$), we choose TSP instances from TSPLIB with node sizes ranging from 52 to 200 and CVRP instances from CVRPLIB with node sizes ranging from 32 to 101 as our test cases.
@@ 5.3 Effectiveness Analysis of ESF + DS
The results presented in Tables 5 (TSPLIB) and 6 (CVRPLIB) indicate that our proposed DS decoder yields results comparable to the well-designed AMDKD and outperforms OMNI-VRP by simply imposing a model architecture improvement method. Remarkably, our method produces solutions identical to the optimal solutions provided by TSPLIB in the following three instances: “pr76”, “pr124”, and “rd100”. To the best of our knowledge, our approach represents the first neural construction model capable of achieving optimal solutions in TSPLIB without requiring iterative update processes. Furthermore, the proposed ESF consistently enhances the generalization performance of the three models. These experimental results further illustrate the effectiveness of enhancing model generalization ability through our proposed components.

### v3

@@ 5.3 Effectiveness Analysis of ESF + DS
The results presented in Tables 5 (TSPLIB) and 6 (CVRPLIB) indicate that our proposed DS decoder yields results comparable to the well-designed AMDKD and outperforms OMNI-VRP by simply imposing a model architecture improvement method. Moreover, as shown in Table 5, the low margin of error indicates that our model exhibits statistically robust performance. Remarkably, our method produces solutions identical to the optimal solutions provided by TSPLIB in the following three instances: “pr76”, “pr124”, and “rd100”. To the best of our knowledge, our approach represents the first neural construction model capable of achieving optimal solutions in TSPLIB without requiring iterative update processes. Furthermore, as shown in Tables 4 and 5, the proposed ESF consistently improves the performance of baseline models and the DS decoder in all test instances with different distribution patterns and shows no preference toward any specific distribution patterns. These experimental results further illustrate the effectiveness of enhancing model generalization ability through our proposed components.
@@ 5.4 Ablation Studies
Furthermore, we assess the robustness of the DS decoder when trained on different distribution patterns. Specifically, we utilize Uniform-rectangle ($\textit{d}_{\textit{Ur}}$), Diagonal ($\textit{d}_{\textit{D}}$), and Gaussian ($\textit{d}_{\textit{G}}$) distribution patterns from [58] for training and compare the performance with those using the DS decoder trained on Uniform ($\textit{d}_{\textit{U}}$), Cluster ($\textit{d}_{\textit{C}}$), and Uniform-Cluster mixed ($\textit{d}_{\textit{M}}$) distribution patterns. As shown in Table 7, the performance differences among the models trained on these diverse distributions are minimal, demonstrating that the DS decoder maintains a high level of robustness when trained on diverse distribution patterns.
Table 7: Results of the DS decoder trained on different distribution patterns
Model, $d_{i}$ TSP-50, optimality gap (%)
$\textit{d}_{\textit{U}}$ $\textit{d}_{\textit{C}}$ $\textit{d}_{\textit{M}}$ $\textit{d}_{\textit{I}}$ $\textit{d}_{\textit{Eo}}$ $\textit{d}_{\textit{Ea}}$
DS, $\textit{d}_{\textit{U}},\textit{d}_{\textit{C}},\textit{d}_{\textit{M}}$ 0.02 0.03 0.04 0.02 0.01 0.03
DS with $n_{\mathcal{D}}=3$ decoders min 0.01 0.02 0.03 0.02 0.01 0.02
max 0.03 0.05 0.06 0.03 0.02 0.05
mean 0.02 0.03 0.04 0.02 0.01 0.03
DS, $\textit{d}_{\textit{Ur}},\textit{d}_{\textit{D}},\textit{d}_{\textit{G}}$ 0.02 0.08 0.07 0.02 0.01 0.09
DS with $n_{\mathcal{D}}=3$ decoders min 0.01 0.05 0.05 0.02 0.01 0.05
max 0.03 0.10 0.10 0.03 0.02 0.11
mean 0.02 0.07 0.07 0.02 0.02 0.08